Science Report April 25th

Hello Mission Support,

Here is the Science Report for 25 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by CGA Eric Shear

1. Emergency EVAC EVA: EVA 6 did not find any suitable locations for evacuation in Lith Canyon. The distance is very far from the hab, difficult to reach and quite exposed to the environment. Although it is beautiful! EVA 7 was to walk south alongside the Hab Ridge Road on foot, but was cut short due to warm conditions. Further investigation of possible locations will be done on future EVAs.

2. Sunspot and Solar Flare Monitoring: The sunspot from the previous observation has split into three smaller ones, and one prominence was seen. For photos, please refer to the astronomy report.

3. Shortwave Texting on EVA’s: After troubleshooting, one Beartooth device continues to not work today. We are awaiting tech support from the manufacturer. In the meantime, we continue to use the note app on our phones to communicate with our deaf crew member.

4. GPS Route Measurements: A GPS tracking route was successfully taken on EVA 7 with the Garmin GPSmap 60Cx. On this EVA, we walked 641 meters (roundtrip).

5. Waterless dish cleaning: Baseline testing was completed this afternoon. 2 swabs were taken from dirty dishes in the sink and two swabs were taken from dishes in the cupboard. The swabs were rubbed on a water/agor mix in petri dishes and will be watched over the next few days for bacteria growth.

Copyright © The Mars Society. All rights reserved. | Main Site